OpenStudy (countrygirl1431):

Write the standard form of the line that is parallel to 2x + 3y = 4 and passes through the point (1, -4). Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

OpenStudy (countrygirl1431):

@jim_thompson5910 @surjithayer

jimthompson5910 (jim_thompson5910):

anything parallel to 2x + 3y = 4 is in the form 2x + 3y = C where C is any fixed number

OpenStudy (anonymous):

any line parallel to ax+by=c is ax+by=k

jimthompson5910 (jim_thompson5910):

plug in (x,y) = (1,-4) into 2x + 3y = C to find the value of C

OpenStudy (countrygirl1431):

how would i do that?

jimthompson5910 (jim_thompson5910):

do you see how x = 1 and y = -4 ?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):